Fermi National Laboratory


Accelerator Update for July 20, 2005 - July 22, 2005

Wednesday July 20
The day shift began with Operations monitoring Tevatron (TeV) store 4285 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with the following colliding, neutrino, and fixed target experiments taking beam: the Collider Detector at Fermilab (CDF), D-Zero (D0), MiniBooNE, NuMI, and the Meson MCenter experiment.

At 8:35 AM, Operations reported that a Linac RF station (KRF2) had tripped off twice with a RF leak indication. Operators reset the station, but at 8:49 AM it tripped off two more times. Linac technicians investigated and fixed the problem.

At 11:29 AM, Operations reported that KRF2 began tripping off again. Linac techs adjusted its leak threshold again.

Operations put the Recycler into shot setup at 12:42 PM to stash antiprotons.

At 1:29 PM, the Main Injector (MI) PSLOOP tripped off on a ground fault indication due a thunderstorm hitting the lab. At the same time the main quadrupole power supplies started running at half of their nominal current. Operators called the Duty Electrician.

The Recycler stash ended at 2:34 PM.

Pbar resumed stacking at 2:38 PM.

At 2:50 PM, MiniBooNE went down to repair DAQ problems. A MADC was replaced at 3:10 PM, but a problem with a power supply (E:H866) held the experiment off. An expert investigated.

Operations established beam to the Meson MTest experiment around 4 PM.

Operations reported at 5:35 PM that experts had repaired E:H866, but that MiniBooNE was still off due to DAQ problems.

MiniBooNE resolved its DAQ problems and resumed taking beam at 7:30 PM.

MTest experimenters reported at 9:23 PM that they were done with the beam for the night.

Thursday July 21
The midnight shift began with Operations monitoring store 4285 and the stash, with Pbar conducting stacking studies, and with MiniBooNE, NuMI, and MCenter.

Operations terminated store 4285 at 12:08 AM.

Operations put the Tevatron (TeV) into shot setup at 12:25 AM.

At 12:40 AM, an Accumulator Septa power supply (A:ISEP2V) failed. An operator called in an expert.

At 1:58 AM, Operations reported a power glitch that caused the TeV bus to trip off. Both the stack and the stash survived. Operators began resetting and rebooting equipment.

At 2:12 AM, the Duty Electrician reported that feeder 47 had tripped off.

Operations established beam to MiniBooNE and NuMI at 3:09 AM.

The Pbar expert reported he had completed his repair on A:ISEP2V at 3:26 AM.

MCenter resumed taking beam at 3:39 AM.

Pbar resumed stacking at 4 AM.

The High Voltage Coordinator reported at 4:10 AM that he was moving D0 to feeder 45 and the Industrial Building Complex to feeder 49.

At 4:12 AM, A:ISEP2V failed again. The expert came back in and changed an interlock card.

The switch to feeder 45 and 49 was completed by 4:55 AM. He also reported that site 38 would be without power.

Pbar resumed stacking at 5:21 AM.

CDF accessed their collision hall at 5:30 AM. They were out of access fifteen minutes later.

Operations put the TeV into shot setup at 6:45 AM.

Operations established combination store 4287 at 8:54 AM with an initial luminosity of 103.49E30. Pbar resumed stacking.

MCenter resumed taking beam at 9:19 AM.

At 9:21 AM, Operations had found the Tevatron Electron Lens (TEL) off and wouldn't reset. Operators investigated and returned it to normal operation.

At 9:24 AM, the TeV suffered an abort at sector A3. There was no quench. Operators found that a sector B0 VCB had opened. Experts began their investigations.

Operators prepared CDF for an access at 10:11 AM.

At 11:37 AM, Booster tripped off due to a LCW pump failure at CUB. Experts switch to a backup pump.

Experts completed the TeV work and CDF came out of access by 2:02 PM.

Operations put the TeV through a dry squeeze at 2:53 PM and then began a study period.

MTest began taking beam sometime around 4 PM.

Operations turned off MiniBooNE at 4:25 PM so they could make repairs.

TeV studies ended at 5:52 PM.

Operations put the TeV into shot setup at 6:39 PM.

MiniBooNE resolved their problem and resumed taking beam at 8:46 PM.

Operations established store 4291 at 10:06 PM with an initial luminosity of 56.63E30.

MCenter and MTest resumed taking beam at 10:16 PM.

Pbar resumed stacking at 10:24 PM.

MTest experimenters said they were finished with the beam at 11:37 PM.

Friday July 22
The midnight shift began with Operations monitoring store 4291, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 12:45 AM, BRF19 tripped off for the first of seven times in about two and a half hours. It reset each time.

At 1:22 AM, Operations reported that a Debuncher quadrupole power supply's (D:Q926) current had dropped. At the same time two other power supplies had their current also drop. Operators called in an expert.

At 2:19 AM, MIRF17 tripped for the first of four times in thirty-five minutes. It reset each time.

The power supply expert reported at 2:32 AM that he had adjusted the shunts causing the D:Q926 current drop.

Operations put the Recycler into shot setup at 2:55 AM to stash antiprotons.

The stash ended and Pbar resumed stacking 3:51 AM.

MCenter resumed taking beam at 3:58 AM.

At 6:47 AM, the TeV suffered a sector E1 quench. Operations investigated and discovered a bulk corrector magnet power supply at F0 had tripped off.

The Plan for Friday and the Weekend
The plans for today are to allow some machine studies during the day and then establish a new store this evening. The plans for this weekend are to stack, stash, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/22/2005   email Fermilab